OpenStudy (jiteshmeghwal9):

If A's income is 20% less than that of B's. How much percent is B's income more than A's ?

OpenStudy (anonymous):

Let x be the income of B

OpenStudy (anonymous):

Then, A's income = x-20% of x

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@jiteshmeghwal9

OpenStudy (anonymous):

A income = 80% B income ??

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

A's income = x-20% of x=0.8x

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (anonymous):

Now, y% of 0.8x = 0.2x

OpenStudy (anonymous):

SOlve for y

OpenStudy (anonymous):

got it?

OpenStudy (jiteshmeghwal9):

sorry ! i didn't gt it :( can u define it to me in an easier way @sauravshakya ?????

OpenStudy (anonymous):

A's income = x-20% of x=0.8x Did u understand up to here

OpenStudy (anonymous):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

nope

OpenStudy (jiteshmeghwal9):

how can A's income=x-20% of x=0.8x ??

OpenStudy (anonymous):

Let B's income be x then... By question, A's income = x-20% of x = 0.8x

OpenStudy (anonymous):

A's income is 20% less than that of B's

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

ohh ! yeahh

OpenStudy (anonymous):

Now, let y be percentage B's income more than A's

OpenStudy (anonymous):

Then, y% of 0.8x = 0.2x

OpenStudy (jiteshmeghwal9):

ok !

OpenStudy (anonymous):

got that?

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (anonymous):

Thus, solve for y

OpenStudy (anonymous):

what did u get?

OpenStudy (jiteshmeghwal9):

ok ! wait a minute

OpenStudy (jiteshmeghwal9):